OpenStudy (anonymous):

Find all y such that the distance between the points (9, -3) and (6, y) is 4

OpenStudy (anonymous):

It's a parabola.

OpenStudy (anonymous):

\[ 4 = \sqrt{(9-6)^2+(-3-y)^2} \]

OpenStudy (anonymous):

\[ 16 = (9-6)^2+(-3-y)^2 \]Solve for \(y\)

OpenStudy (anonymous):

distance between the points we do this : \[\sqrt{(x1-x2)^2+(y1-y2)^2}\]

OpenStudy (anonymous):

No matter what i do it still comes up with the wrong answer. Can you show me hoe to break it down more?

OpenStudy (anonymous):

what ????

OpenStudy (anonymous):

when i solve for x it never is the right answer

OpenStudy (anonymous):

heh !! u think ! Please find ! I m there to help u . And if it was fall there s no problem because our answers are not always right !:) Now I m there to help not answer ! Now start with full power !:)